Snow Removal - By 1PM, Shani had already shoveled 2 driveways, and by 6PM that day, the number was up to 7.
A) Find Shani's average shoveling rate, in number of driveways per hour.
B) Find Shani's shoveling rate,in hours per driveway.

Answer by stanbon(75887)

Snow Removal - By 1PM, Shani had already shoveled 2 driveways, and by 6PM that day, the number was up to 7.
-----
You have 2 points:: (0,2), (5,7)
slope = (7-2)/(5-0) = 5/5 = 1
---------
A) Find Shani's average shoveling rate, in number of driveways per hour.
Ans: average = slope = 1 driveway per hour
----------
B) Find Shani's shoveling rate,in hours per driveway.
Ans: rate = 1 hr per driveway
-----------------------------
Cheers,
Stan H.